OpenStudy (anonymous):

Find the LCM of 25y^2+100y and 5y^2+25y+20

OpenStudy (anonymous):

25y^2+100y=25y(y+4) 5y^2+25y+20=(5y+5)(y+4) so LCM=25y(y+4)(5y+5)

OpenStudy (anonymous):

the answer was 25y(y+4)(y+1) this problem really stumps me....

OpenStudy (mathmate):

25y^2+100y and 5y^2+25y+20 25y^2+100y = 25y(y+4) 5y^2+25y+20=5(y^2+5y+4)=5(y+4)(y+1) To find the LCM, find the product of the factors, making sure common factors are not repeated, thus LCM=25y(y+4)(y+1)